An electronics company makes two types of switches. Type A takes 4 minutes to make and requires $3 worth of materials. Type B takes 5 minutes to make and requires $5 of materials. In the latest productionbatch, it took 32 hours to make these switches, and the materials cost $1740. How many of each type of switch were made?
Answer by ikleyn(53763) 
 .
4a + 5b = 32*60 (1) (minutes to make)
3a + 5b = 1740 (2) (material cost)
To solve the system, use the Elimination method. For it, subtract eq(2) from eq(1). You will get
4a - 3a = 32*60 - 1740 = 180.
a = 180.
Then from eq(2), 5b = 1740 - 3*180 = 1200, b = 1200/5 = 240.
ANSWER. 180 switches of the type A and 240 switches of the type B.
|
|
|